A bank currently has $70,000 in transaction deposits, $6,000 in cash in the vault, $12,000 on deposit with the Fed, and $7,000 in government securities. The required reserve ratio is 20 percent. Answer these questions:

1) What is the maximum amount the money supply can increase, assuming this bank is the only bank in the system that has excess reserves? You need to know how to calculate actual reserves, required reserves, excess reserves, and the money multiplier to solve this problem. All these definitions are in your text, on or about page page 335 in your text.

2) An individual deposits a $750,000 check into the bank. That individual had just converted foreign currency into dollars so the $750,000 was not in the money supply before the deposit. What is the maximum size loan the bank can make once the check clears?

You must post an initial post and make at least two comments on your classmates' posts.
Answer #1

1.

Required Reserve ratio = 20%

So, required reserve = 70000*20% = $14000

Actual reserve = (6000+12000) = $18000

Excess reserve = (6000+12000) - 14000 = $4000

Money multiplier = 1/20% = 5

Hence,

Maximum amount of money supply to be created = 4000* money multiplier = 4000*(1/20%)

Maximum amount of money supply to be created = $20000

2.

Size of the deposit = $750000

Maximum size of the loan = 750000 - 750000*20%

Maximum size of the loan = $600000